### v1

# Prime and Reach: Synthesising Body Motion for Gaze-Primed Object Reach

Abstract
Human motion generation is a challenging task that aims to create realistic motion imitating natural human behaviour. We focus on the well-studied behaviour of priming an object/location for pick up or put down – that is, the spotting of an object/location from a distance, known as gaze priming, followed by the motion of approaching and reaching the target location. To that end, we curate, for the first time, 23.7K gaze-primed human motion sequences for reaching target object locations from five publicly available datasets, i.e., HD-EPIC, MoGaze, HOT3D, ADT, and GIMO.
We pre-train a text-conditioned diffusion-based motion generation model, then fine-tune it conditioned on goal pose or location, on our curated sequences. Importantly, we evaluate the ability of the generated motion to imitate natural human movement through several metrics, including the ‘Reach Success’ and a newly introduced ‘Prime Success’ metric. On the largest dataset, HD-EPIC (LABEL:fig:egoallo_examples), our model achieves 60% prime success and 89% reach success when conditioned on the goal object location.
1 Introduction
Text-guided human motion generation has seen significant advancements, generating increasingly realistic body motions [tevet2023mdm, guo2023momask, Guo2022t2m]. These works excelled at translating textual descriptions into a wide array of human movements, ranging from simple actions like walking and running to more complex motions such as dancing. The scope has then expanded to include navigating virtual environments towards a spatial target location [diomataris2024wandr, karunratanakul2023gmd, xie2024omnicontrol]. More recently, models are designed to generate human motions of interactions with various objects [taheri2022goal, paschalidis20253d, tendulkar2023flex, xu2023interdiff, diller2024cg], such as grasping small objects or interacting with furniture. However, these works heavily rely on synthetic datasets [tendulkar2023flex, araujo2023circle] or datasets collected in lab settings [taheri2020grab, jiang2023full, lv2024himo]. As a result, they do not capture or model the prime and reach (P&R) behaviour, synthesising motions that are less natural and thus less useful as a digital replica of human behaviour.
On the other hand, egocentric data collection provides a compelling alternative solution, enabling the capture of more natural interactions directly from the first-person perspective, accompanied by a gaze-mounted camera. Several datasets capture natural interactions in diverse daily activities [ma2024nymeria, grauman2024ego] and human-human interactions [zhang2022egobody], as observed from an egocentric perspective. Nevertheless, most egocentric datasets [Damen2022RESCALING, perrett2025HD-EPIC, grauman2022ego4d] do not capture full-body motion, as pairing egocentric and 3D sensors is costly and challenging. Several recent works [li2023ego, yi2025egoallo, guzov-jiang2025hmd2, chi2025estimating] have explored estimating full-body motion conditioned on the egocentric camera pose and viewpoint, training on diverse data [ma2024nymeria, grauman2024ego]. Although all these egocentric datasets capture diverse human behaviour, they have not been explored for modelling the preparatory priming and reaching motion (see Fig LABEL:fig:egoallo_examples for sample sequences).
Human visuomotor coordination is fundamentally anticipatory, continuously integrating sensory information to facilitate fluid, goal-directed actions [jia2025learning]. The anticipatory mechanism heavily relies on gaze. By directing visual attention to relevant objects or areas before reaching, gaze offers crucial predictive signals that enable the motor system to prepare and execute actions efficiently and naturally [land1999roles, johansson2001eye, hayhoe2005eye, jia2025learning]. The ability to prime and reach has recently been investigated in various real-world applications, including robotics [kerrj2025eyerobot, rahrakhshan2022learning, shafti2019gaze]. For example, in a manipulation involving a towel, the robot would visually fixate on the towel and the grasp point before extending its arm [kerrj2025eyerobot]. However, such behaviour has not been explored in training or evaluating motion generation.
In this work, we enable the synthesis (or generation) of priming and reaching behaviour by combining egocentric datasets that offer gaze priming and full-body motion with conditioned diffusion models.
We list our contributions as follows:
  • We explore, for the first time, synthesising full-body motion to replicate humans’ natural ability to first gaze-prime then approach and reach objects for interactions.
  • We curate five datasets, totalling 23.7K prime and reach sequences from five public datasets.
  • We train goal-conditioned motion generation models on our curated datasets and evaluate their ability to generate prime and reach motion. We compare conditioning on the goal pose vs the goal object location.
  • We introduce the ‘Prime Success’ metric to particularly evaluate the ability to synthesise gaze priming behaviour.
  • Our results show that using goal pose as a condition, our model can boost priming ability by up to 1818x compared to previous methods, and can similarly improve reach success by up to 88x. When conditioning on the target object location only, our model improves the priming ability by up to 55x.
2 Related Work
2.1 Human Motion Generation
Human motion generation aims to create realistic, continuous human movements that simulate or animate natural human motion. The majority of work in this field generates human motion from a single modality such as text [Guo2022t2m, petrovich22temos, petrovich23tmr, zhong2023attt2m, athanasiou2022teach], action [guo2020action2motion, cervantes2022implicitneuralrepresentationsvariable, petrovich21actor, posegpt], speech [zhu2023taming, alexanderson2023listen], and music [li2021learn, tseng2022edge, sun2022nonfreezing], with a few recent works tackle motion generation from multiple modalities [li2024unimotion, genmo2025, bian2024motioncraft, luo2024m3gpt].
Text-to-motion Generation.
Initial efforts [ahuja2019language2pose, ghosh2021synthesis] in text-to-motion were deterministic, converging to an averaged motion given an input text. After the advent of the denoising diffusion models [ho2020denoising, song2021denoising], these models are nowadays a common practice to generate text-conditioned human motion [tevet2023mdm, zhang2024motiondiffuse, zhang2023remodiffuse, kim2023flame]. MDM [tevet2023mdm], MotionDiffuse [zhang2024motiondiffuse], and FLAME [kim2023flame] are conditioned on features extracted by a pre-trained text encoder. Another common approach is to disentangle motion representation from generation [guo2023momask, jiang2024motiongpt, chuan2022tm2t, zhang2023t2mgpt]. This is a two-stage process: first, a VQ-VAE [van2017neural] is trained to create a discrete codebook that tokenises continuous motion sequences. Then, a Transformer-based autoregressive model is trained on these discrete tokens to learn the motion primitives by predicting the next token in a sequence. Although the text can serve as a strong signal for conditioning semantic motion, these methods often lack precise control over body positions.
Location-Conditioned Motion Generation.
Our work is related to a line of research that generates human motion conditioned on a target location. Several works extend the text-to-motion model to enable various spatial controls (e.g., entire root trajectory [karunratanakul2023gmd], keyframe root location [karunratanakul2023gmd], obstacle avoidance [karunratanakul2023gmd], temporally and spatially sparse joints [pinyoanuntapong2025maskcontrol, xie2024omnicontrol], or goal location [tevet2025closd]). Recently, WANDR [diomataris2024wandr] introduced a data-driven model conditioned on the initial pose and the goal location of the right wrist to generate avatars that walk and reach the goal in 3D space. Other works [li2024egogen, diomataris2025movinglookingvisiondrivenavatar] address the goal-reaching human motion generation via egocentric perception and reinforcement learning. Despite these advancements, these methods rely on synthetic datasets [araujo2023circle] or MoCap-based datasets [mahmood2019amass], which limit their ability to generate natural interactions in real-world scenarios. Additionally, these works do not address or evaluate priming. In this work, we curate the first set of datasets that include full-body, priming, and reaching, with a focus on replicating this human priming-then-reaching motion through generation.
Ego-body Pose Generation.
Recent research has explored estimating [jiang2022avatarposer, jiang2024egoposer, Luo2021DynamicsRegulatedKP, chi2025estimating, guzov-jiang2025hmd2] or forecasting [yuan2019ego, escobar2025egocast, patel25uniegomotion, Hatano2025EgoH4] human motion from an egocentric perspective. These methods typically adopt a generative approach as the human body is largely invisible from an egocentric view, unlike ego-body pose estimation from a downward-facing camera [Millerdurai_EventEgo3D_2024, wang2024egocentric, hakada2022unrealego, tome2019xr, cuevas2024simpleego, zhao2021egoglass]. EgoEgo [li2023ego] is the first work to propose head pose (camera pose) conditioned human motion generation, but was mainly evaluated on synthetic datasets. Subsequently, EgoAllo [yi2025egoallo] proposes a head-centric representation (i.e., canonicalisation) to achieve spatial and temporal invariance, and also enables the integration of in-view 3D hand poses for better prediction. We utilise the ego-body pose estimation method [yi2025egoallo] to generate human motion on gaze-primed and reach sequences curated from egocentric datasets.
2.2 Eye-gaze in Motion
Eye-gaze is an important predictive signal that directs attention and primes the processing of future movements [land1999roles, johansson2001eye, hayhoe2005eye]. Recognising the critical role of gaze, recent research has been focusing on estimating gaze/saliency [lai2022eye, lai2024eye, chong2018connecting] or explicitly leveraging this cue for various problems, such as video understanding [mazzamuto2025gazing, peng2025eyemllmbenchmarkingegocentric] or human-robot interactions [stolzenwald2018can, saran2018human]. Several works focus on future motion prediction following gaze priming [wei2018and, yan2023gazemodiff, hu2025hoigaze, lou2024multimodal, jia2025learning]. Tian et al. [tian2024gaze] generate hand-object interactions but only in table-top settings. Different from these works, we wish to synthesise both the gaze priming and the reach motion, for the full body, conditioned on the goal.
3 Prime and Reach Data Curation
We first introduce the principle of curating ‘Prime and Reach’ (P&R) sequences from longer videos. We then detail the steps we carried out to curate these sequences from five public datasets. We note the statistics of these sequences, which we use for training and evaluation.
3.1 P&R sequence Curation
Interaction datasets include multiple and frequent object reaching and manipulation behaviours. However, a critical aspect largely unexplored is the role of gaze in priming or “spotting” objects prior to the reaching motion. We take this missed opportunity and curate for the first time P&R motion sequences from datasets capturing wearable gaze and object interactions. We are inspired by the “gaze priming” discussed in [perrett2025HD-EPIC] where objects were annotated in 3D and then used to identify the fixation that occurs prior to the physical action, signalling intent of interaction.
Starting from long videos, we extract timestamps for object pick-up or put-down events. We note that priming takes place also during put-down where the future location of an object is primed before the action. Given the known pick/put event at time $t_{e}$, we analyse a temporal window of duration $w$ immediately preceding it to find a moment $t_{p}\in[t_{e}-w,t_{e}]$. We wish to identify when the user’s gaze first attends to or primes the relevant location for the pick-up/put-down event. For pick-up events, we associate the event with 3D location of the object. This location will be used to identify the priming event. Importantly, for put-down events, we instead use the future 3D location of the object (which at the start of the motion is an empty part of the 3D space) to search for the priming – i.e., we track the intersection of the gaze of the camera wearer with this empty space, priming the location where the object is going to be placed.
Specifically, we project the user’s gaze into the 3D environment to form a ray. First, the gaze direction provided by the eye-tracker in the camera’s local coordinate system at any time $t$, $\mathbf{p}^{t}_{\text{gaze\_cam}}$, is transformed using the camera-to-world transformation matrix $\mathbf{T}^{t}_{\text{c2w}}$. The final normalised gaze direction vector, $\hat{\mathbf{d}}^{t}_{\text{gaze}}$, is then computed as the vector from the camera’s world position, $\mathbf{o}^{t}_{\text{cam}}$, to this new world-space gaze point $\mathbf{p}^{t}_{\text{gaze\_world}}$ as shown in Equation 1.
$\mathbf{p}^{t}_{\text{gaze\_world}}=(\mathbf{T}^{t}_{\text{c2w}}\mathbf{p}^{t}_{\text{gaze\_cam}})$
[…]
We register a relevant location as primed if the gaze ray, originating from the camera’s position, intersects with the corresponding 3D bounding box or 3D location $o_{\text{3D}}$. Therefore, we define the prime time $t_{p}$ as
$T_{\text{int}}=\{t|t\in[t_{e}-w,t_{e}],\mathbb{I}(\text{intersect}(\hat{\mathbf{d}}^{t}_{\text{gaze}},o_{\text{3D}}))=1\}$
$t_{p}=\min_{t\in T_{\text{int}}}t,$ (2)
where $T_{\text{int}}$ is the set of all timestamps within the temporal window $t_{e}-w,t_{e}]$, where the gaze ray intersects the 3D location and $t_{p}$ is the first moment where the intersection happens. We discard sequences where $T_{\text{int}}=\emptyset$. Following [perrett2025HD-EPIC], we use $w=10$ secs. To compute the intersection i.e. $\text{intersect}(\hat{\mathbf{d}}^{t}_{\text{gaze}},o_{\text{3D}})$, we use the slab test method [Majercik2018Voxel], details of which are available in the supplementary. At the end of this process, we get P&R sequences each defined by a prime time $t_{p}$ and reach time $t_{e}$.
Table 1: Curated Dataset Statistics. We report statistics on curated P&R sequences across five publicly available datasets, ordering them by the size of curated sequences. We report the number of P&R sequences, duration between prime time and reach time i.e. $t_{e}-t_{r}$ (Prime Gap), body pose type, the distance/movement of body and hand. * indicates that body poses are estimated using [yi2025egoallo].
Dataset #P&R Seq.
Sequence
Duration ($s$)
Prime
Gap ($s$)
Body Pose Type
Body
Movement ($m$)
Hand
Movement ($m$)
HD-EPIC [perrett2025HD-EPIC] 18,134 $5.49\pm 2.76$ $3.55\pm 2.79$ SMPL-H* $0.72\pm 0.67$ $0.45\pm 0.22$
MoGaze [kratzer2020mogaze] 2,637 $3.64\pm 0.94$ $1.53\pm 0.92$ 3D Skeleton $1.07\pm 0.62$ $0.75\pm 0.25$
HOT3D [banerjee2025hot3d] 2,416 $4.31\pm 1.54$ $2.37\pm 1.57$ SMPL-H* $0.20\pm 0.15$ $0.38\pm 0.19$
ADT [pan2023aria] 411 $7.44\pm 2.47$ $4.51\pm 2.74$ 3D Skeleton $1.23\pm 1.28$ $0.56\pm 0.24$
GIMO [zheng2022gimo] 130 $7.11\pm 2.49$ $4.47\pm 1.49$ SMPL-X [pavlakos2019expressive] $3.09\pm 1.20$ $0.64\pm 0.19$
Figure 2: Examples of curated P&R motion sequences from five different datasets.
3.2 Datasets
As explained in Sec. 3.1, we formulate how P&R sequences can be curated from long video sequences. We consider five publicly available human-object interaction datasets, all of which contain 2D gaze tracked from wearable gaze trackers along with camera poses [perrett2025HD-EPIC, kratzer2020mogaze, banerjee2025hot3d, pan2023aria, zheng2022gimo].
[…]
MoGaze [kratzer2020mogaze] is human motion data designed explicitly for human-object interactions, with a particular focus on ‘pick’ and ‘put’. The dataset includes synchronised full-body motion captured using motion capture markers, 3D object models, and 6-DoF and eye-gaze data. The dataset contains 180 minutes of motion capture data from seven participants performing pick-and-put actions along with temporal segment annotations of these actions. Using the gaze data and the object locations, we determine priming timestamps ($t_{p}$) for each pick and put event. We slice the motion data for $t\in[t_{p}-2,t_{e}]$ constructing $2,637$ P&R sequences.
HOT3D [banerjee2025hot3d] captures 3D hand-object interactions. The dataset offers 198 ARIA recordings featuring 14 subjects interacting with 33 diverse objects. We only use the ARIA videos as these provide gaze information. As pick/put timestamp annotations are not provided, we extract temporal segments where an object is in-hand by thresholding (<5<5 cm) the distance between the nearest hand vertices and the object locations. After identifying these in-hand segments, we refine the segment boundaries by detecting the object state change from stationary to in-hand or vice-versa. This gives us pick/put events $t_{e}$. We use gaze and camera pose to estimate the prime time for these events, resulting in $2,416$ P&R sequences. We estimate full-body motion for these sequences using EgoAllo [yi2025egoallo].
Aria Digital Twin (ADT) [pan2023aria] provides a rich collection of synchronised data, including images, eye-tracking data, 6-DoF object data, and 3D human poses. 72 videos in the dataset capture indoor activities and interactions involving 398 unique objects and provide paired eye gaze and 3D body motion data. We curate P&R sequences from these videos. Same as HOT3D, we find temporal segments when objects are in-hand by thresholding the distance between object locations and nearest wrist locations from the corresponding body poses. We identify pick and put events near the temporal boundaries of these segments based on how the object state changed. These events were then primed to determine $t_{p}$, resulting in 411 P&R sequences with full-body motion data.
GIMO [zheng2022gimo] is a benchmark that focuses on intent-guided human motion segments. It provides 217 trimmed segments, along with corresponding SMPL-X fitted IMU-captured body poses and egocentric views with eye gaze data captured by the HoloLens 2. We discard all segments that include resting activities e.g., sitting or lying on the bed. Gimo segments include object pick-up but no put-down. However, they do not provide the 3D locations or timestamps of object pick-up. Therefore, we manually annotate these event timestamps ($t_{e}$) from RGB videos and use the relevant wrist location at the timestamp as our object locations. We determine $t_{p}$ for each of these events following the same method as before. We get 130 P&R SMPL-X body motion sequences.
In total, we curate $23,728$ P&R sequences from the five datasets. Statistics are provided in Tab. 1. We showcase sample P&R sequences in Fig. 2. Importantly, we unify body pose formats across datasets by representing them using the canonicalised 22-joint body motion used in HumanML3D [Guo_2022_CVPR]. Following [tevet2023mdm], we convert the 22 joint positions to 263-dim vector representation that combines local pose, rotation and velocity of each joint. The curated sequences from each dataset are split 70%-30% into train and test sets. Details are provided in supplementary.
4 Method
4.1 Objective
Here, we address the task of goal-conditioned human motion generation with the ability to prime and reach a given object. Specifically, the task aims to generate human motion sequences $\{x^{i}\}_{i=1}^{N}$ of length $N$, where $x^{i}\in R^{J\times 3}$ represents 3D positions of $J$ body joints, guided either by desired goal pose or target goal object location as a condition. We consider and compare the two conditions. As sequences can be of varying length, we opt to predict a fixed-length sequence by uniformly up-/down-sampling $N$ frames. Predicted sequences can then be up- or down-sampled temporally so that the motion length is the same as the ground truth.
4.2 Prime & Reach Motion Diffusion Model
Conditioning. Diffusion models have demonstrated exceptional capability for text-conditioned motion generation [tevet2023mdm, zhang2023remodiffuse]. Motivated by this, we use a diffusion generative model, as in [tevet2023mdm] for our task. We present our architecture in Fig. 3. Starting from pure noise at $t=T$, the transformer decoder generates motion through iterative denoising over multiple diffusion timesteps $t=\{T,...,0\}$ where $t=0$ produces the predicted motion. This generation is guided through a set of conditions injected into the decoder. We condition our prime and reach motion generation on –
  • Text prompt: this allows the model to benefit from text-to-motion pre-training. We describe the action as e.g., ‘The person moves across and picks/puts an object.’ We use the knowledge of the action (i.e., whether it is a pick up or a put down) in both training and inference to guide the synthesis. We refer to this conditioning text as $\mathbf{c}$.
  • Initial state of the body describing where and how the motion initiates. As P&R sequences do not start from a static or neutral pose, but are sampled from within a longer sequence, it is important to feed the initial pose and the velocity at the first frame, as this impacts the guided motion. We represent this by the starting pose $(\hat{x}^{1})\in R^{J\times 3}$ and the joint velocities i.e. $(\hat{x}^{1}-\hat{x}^{0})$ where $\hat{x}^{0}$ is the preceding frame before the curated P&Rsequence. We flatten and concatenate the start pose and velocities into a single vector, which forms our initial state.
  • Goal. We evaluate two possible goal formulations. The first is the complete goal/target pose at the end of the motion $(\hat{x}^{N})$. The goal pose not only guides the motion to reach the object but additionally guides where the agent would stand relative to the object (through the full pose at the end of the motion) as well as which hand would be reaching the object (guided through the position of the hand joints). Second, we use the more challenging goal of only specifying the object location $(o_{3D}\in R^{3})$ as a condition. For this goal, the model needs to estimate where the body should be relative to the object and which hand to be the one reaching for the object.
Architecture. Next, we explain how we use these conditions for the generation process. First, we encode the input text prompt $\mathbf{c}$ using a pre-trained text encoder [sanh2019distilbert]. Both the diffusion noise time step $t$ and $\mathbf{c}$ are then projected to a latent space and summed to get a token $\mathbf{z_{t}}$. Note that for generating motion conditioned on text, $\mathbf{z_{t}}$ is directly injected into the transformer decoder’s layers. We pre-train this model for text-conditioned motion generation. This pre-training enables the model to learn prior knowledge of fine-grained full-body motion involved in everyday activities.
For learning P&R motion generation, we initialise our model with the pre-trained text-to-motion conditioned weights and fine-tune it with all three conditions. To add the initial state and goal condition, we first flatten and concatenate them into a single 1D vector. The resultant vector is linearly projected to a latent space to give $\mathbf{p}$. $\mathbf{z_{t}}$ is modified by adding condition $\mathbf{p}$ to $\mathbf{z_{t}}$.
$\mathbf{\tilde{z}_{t}}=\mathbf{z_{t}}+\mathbf{p}$ (3)
This modified condition $\mathbf{\tilde{z}_{t}}$ is now injected into the transformer decoder’s layers through cross-attention blocks. This guides the motion sequence generation over multiple diffusion timesteps. We ablate other options of injecting p, including cross-attention, in the supplementary.
Once de-noised, the decoder produces 263-dim representations of body joints $\{v^{n}\}_{n=1}^{N}$ where $v^{n}\in R^{263}$ combines local position, local rotation and local velocity of all 22 body joints. This is post-processed as $\{x^{n}\}_{n=1}^{N}=g(\{v^{n}\}_{n=1}^{N})$ to get the predicted 22 joint positions. We use the same $g$ as in [karunratanakul2023gmd, tevet2023mdm].
[…]
Figure 3: P&R motion diffusion model for goal-conditioned motion generation. We concatenate the initial state of the human body and the goal pose/goal object as conditions, along with a text condition describing the type of action the motion is expected to perform. This accumulated condition is injected into the transformer decoder layers, which then outputs an NN-length motion sequence over multiple diffusion steps.
5 Experiments
Here we explain implementation details (Sec. 5.1), evaluation metrics (Sec. 5.2), baselines (Sec. 5.3), experimental results (Sec. 5.4) and ablations (Sec. 5.5).
5.1 Implementation Details
We pre-train our P&R motion diffusion model for text-conditioned motion generation on the large-scale Nymeria [ma2024nymeria] dataset to learn the motion prior of everyday activities. Nymeria provides large-scale full-body motion data of participants performing diverse actions captured by Xsens mocap sensors, accompanied by atomic narrations describing the actions. The narrations are used as text guidance. For pre-training, an initial learning rate of 1e41e-4 is used, for a maximum of 600K steps. We use a motion length of $N=150$ and classifier-free guidance with a probability of 0.2. The pre-training takes $\sim 36$ hours on one H200 GPU.
Initialised with the pre-trained weights, we fine-tune our Prime and Reach model on the curated P&R sequences from each dataset separately. This is because various datasets capture different types of activities. For fine-tuning, we use a learning rate of 5e55e-5 for 250K steps, taking around 15-20 hours on HD-EPIC and HOT3D. For MoGaze, ADT and GIMO, we use a learning rate of 1e51e-5 for $\sim 300K$, $\sim 150K$, and $\sim 170K$ steps respectively. We use $T=50$ diffusion steps in pre-training, finetuning, and inference, following [tevet2023mdm].
5.2 Evaluation Metrics
We report results on six metrics: two to directly evaluate our ability to prime and reach, two to evaluate the body pose at the goal, and two to evaluate the entire generated motion.
[…]
$\frac{1}{B}\sum_{b=1}^{B}\mathbb{I}\{\min_{t=t_{p}-\sigma}^{t_{p}+\sigma}(\cos^{-1}(\mathbf{\hat{H}_{t_{p}}},\mathbf{H_{t}}))\leq\theta\},$ (4)
where $B$ is the total number of motion sequences, $\mathbb{I}$ is the indicator function and $\sigma$, $\theta$ are hyperparameters. Unless mentioned otherwise, we use $\theta=16\deg$ and $\sigma=0.2$ sec. Prime Success gives the percentage of sequences where this angular error is $\leq$ threshold $\theta$. Having $\sigma$ relaxes the metric to allow predicted motion to prime the object in a temporal window of $2\sigma$ around $t_{p}$.
(2) Reach Success. We follow [diomataris2024wandr] and evaluate our reaching ability. This calculates the percentage of sequences where either wrist in the predicted motion reaches within 10 cm of the goal at $N$.
(3) Location Error. Also used in [karunratanakul2023gmd], we calculate the distance between the final pelvis locations of the ground truth and predicted motions and report the percentage of sequences where the distance is $\geq$ (50 cm).
(4) Goal MPJPE, which calculates the error between the final full body pose of the ground truth and the predicted motion. This includes the error in all joints including the hand reaching the object. Notice that this error assumes the same hand is reaching out to the object as the ground truth.
(5) Mean Per Joint Position Error (MPJPE), which averages the joint position error in Euclidean distance over all generation frames.
(6) Foot Skating, a common evaluation of the generated human motion [karunratanakul2023gmd], which measures the proportion of frames where either foot slides with velocity (>50>50 cm/s) between consecutive frames while maintaining contact with the ground (foot height <5<5 cm).
Table 2: Comparison of motion generation baselines on our curated P&R sequences using different metrics. We show results for HD-EPIC, MoGaze, HOT3D, ADT and GIMO separately. The baselines are grouped based on the kind of condition they use for generation. $\dagger$ denotes that the model is trained on HumanML3D. denotes the model is trained on Nymeria.[ma2024nymeria].
HD-EPIC MoGaze
Condition Method
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
No condition Static 30.15 15.02 0.82 50.92 0.45 1.64 4.17 1.03 74.26 0.64
Text MDM $\dagger$ 4.33 0.50 0.94 71.52 0.66 0.21 10.19 3.20 2.03 93.01 1.45 0.29
MDM 5.82 1.66 0.83 58.82 0.58 0.03 1.86 1.79 1.18 78.72 0.74 0.07
MDM 30.52 18.76 0.84 58.73 0.54 0.04 14.43 7.74 1.24 85.42 0.72 0.27
+ Initial State & Goal Pose GMD $\dagger$[karunratanakul2023gmd] 29.23 26.30 0.30 4.00 0.32 0.17 2.01 50.24 0.35 2.23 0.54 0.11
P&R 66.31 91.23 0.09 0.20 0.17 0.13 35.34 82.96 0.09 0 0.24 0.21
+ Initial State & Object Loc. WANDR $\dagger$[diomataris2024wandr] 10.21 95.66 0.56 54.85 0.58 0.21 11.26 75.62 0.60 64.91 0.56 0.30
P&R 59.06 89.48 0.30 16.28 0.24 0.12 48.51 92.78 0.39 22.32 0.33 0.25
HOT3D ADT
Condition Method
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
No condition Static 42.13 31.38 0.30 6.96 0.27 7.29 11.46 1.98 72.92 1.12
Text MDM $\dagger$ 19.10 2.45 0.49 38.01 0.44 0.10 14.58 3.65 2.72 94.79 1.85 0.24
MDM 0.89 0.35 0.54 42.34 0.46 0.01 7.81 6.25 2.15 89.06 1.25 0.07
MDM 36.44 30.32 0.36 14.45 0.32 0.00 8.85 8.85 2.23 90.63 1.23 0.17
+ Initial State & Goal Pose GMD $\dagger$[karunratanakul2023gmd] 34.04 11.70 0.41 28.73 0.36 0.02 20.69 55.17 0.42 15.79 0.48 0.14
P&R 75.26 90.69 0.12 0.53 0.11 0 28.65 67.71 0.22 7.81 0.31 0.18
+ Initial State & Object Loc. WANDR $\dagger$[diomataris2024wandr] 12.26 90.41 0.58 59.70 0.52 0.07 14.94 85.71 0.73 79.87 0.66 0.32
P&R 67.50 94.50 0.16 2.35 0.12 0 25.00 75.52 0.56 53.15 0.47 0.22
GIMO
Condition Method
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Goal
MPJPE
$\downarrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Foot
Skating
$\downarrow$
No condition Static 14.28 0 2.85 100 1.84
Text MDM $\dagger$ 0 0 5.26 100 2.23 0.17
MDM 4.76 0 3.95 100 2.10 0.04
MDM 0 0 2.96 100 1.64 0.14
+ Initial State & Goal Pose GMD $\dagger$[karunratanakul2023gmd] 0 4.76 0.44 11.90 0.62 0.43
P&R 42.12 38.12 0.41 15.32 0.54 0.15
+ Initial State & Object Loc. WANDR $\dagger$[diomataris2024wandr] 19.04 71.42 0.75 85.71 0.77 0.21
P&R 47.61 71.42 0.54 52.38 0.62 0.16
Figure 4: P&R performance for pick v/s put.
5.3 Baselines
To assess the ability of generated motion to replicate humans’ P&R behaviour, we benchmark five methods on our curated datasets: one naive baseline and three previous works (one from the text-to-motion generation and two from location-conditioned human motion generation):
  • Static is a naive baseline that uses the average full-body pose of training data and keeps it static for the entire motion sequence. It showcases the difficulty of the dataset.
  • MDM [tevet2023mdm]. We evaluate the checkpoint trained on HumanML3D [Guo_2022_CVPR] vs our pre-training on Nymeria [ma2024nymeria]. We also fine-tune this model, pre-trained on Nymeria, on each dataset using only text conditioning.
  • GMD [karunratanakul2023gmd], a guided motion diffusion trained on HumanML3D based on UNet architecture, is a two-stage process designed for controllable human motion synthesis with text. The first stage generates a root trajectory that guides full-body motion generation in the second stage.
  • WANDR [diomataris2024wandr] is a data-driven, autoregressive method that uses a conditional Variational Auto-Encoder (c-VAE) to generate motion frame-by-frame, trained on the combined dataset of AMASS [mahmood2019amass] and CIRCLE [araujo2023circle]. The model is guided by intention features that encode the goal’s position and the time remaining to reach it.
5.4 Results
In Tab. 2, we compare our proposed P&R diffusion model against baselines on five different datasets. The naive static baseline performs poorly on all the metrics for all datasets. Its poor prime success ($20\%$ on average) and reach success ($12\%$ on average) highlight the difficulty of the task on our curated datasets.
Text-conditioned baselines rely solely on the knowledge of the pick/put action to generate motion and have no information about the target location to prime or reach. We report the performance of three variants of MDM [tevet2023mdm]. These baselines perform poorly on all the metrics as they lacks sufficient guidance. Finetuning MDM on the target datasets helps generate better prime and reach motions with a maximum gain of $24.7\%$ in prime success and $17.1\%$ in reach success on HD-EPIC. The fine-tuned MDM only marginally improves over the static baseline (average of $20.0\%$ prime success and $13.1\%$ reach success), showing that text is not a sufficient condition for prime and reach.
[…]
Figure 5: Qualitative results on 3 datasets: Ground truth sequence in light green, goal-pose conditioned prediction in translucent yellow, and target location conditioned generation in brown. We show the pose at the initial, prime, and reach timesteps. Prime direction for both ground truth and predictions are shown using arrows, and target object location is shown in sphere.
Conditioning with Object Location. WANDR [diomataris2024wandr] achieves strong success scores over all datasets, as that the model’s reward focuses on the reaching of goal. However, it underperforms on all other metrics, including the metrics that evaluate the motion between the initial state and target, and it is very poor at priming, with an average of only $16\%$ over all datasets. P&R trained using initial state and object location as condition achieves significant improvements in prime success with gains ranging from $11.9\%$ on ADT to $48.8\%$ on HD-EPIC. It reaches or outperforms WANDR in reach success on three out of the five datasets. It also outperforms WANDR by a big margin on all other metrics.
Importantly, when comparing the object location condition with the goal pose, we note that metrics that evaluate the goal location (Goal MPJPE and Loc Error) surely benefit from the guidance on where the person is exactly standing at the end. However, when focusing on the prime and reach success, object location can outperform the stronger goal pose on MoGaze and GIMO. We find that due to limited training data, performance on small-scaled ADT and GIMO are evidently worse.
Results for pick and put. We analyse the performance of the P&R model separated by the action (i.e. pick up or put down) in Fig. 4. P&R performs well on both, but pick-up motions are relatively more challenging than put-down actions, especially for priming ability on MoGaze.
Qualitative Results. We demonstrate qualitative examples of our P&R generated motions in Fig. 5. Generated P&R motions appear natural. Starting with an initial pose and velocity, our generated motion first primes the target object (see arrow) and then reaches to it with one of the hands. Evidently, using the goal location matches better the ground-truth. However, using the object location condition solely successfully synthesises reach but positions the body in a different location at the goal. We demonstrate this ability on various target locations including challenging ones, where the location can be located low or high.
Table 3: Impact of condition: We show how each of our modified condition impact P&R model’s performance (last row).
HD-EPIC MoGaze
Object
Loc.
Initial
Pose
Initial
Vel.
Text
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc
Err
$\downarrow$
MPJPE $\downarrow$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc
Err
$\downarrow$
MPJPE $\downarrow$
40.14 30.62 45.50 0.38 18.45 10.94 73.21 0.49
38.71 83.44 27.30 0.37 41.07 87.05 28.57 0.46
42.84 88.15 22.41 0.34 36.83 87.57 24.33 0.35
56.76 86.31 16.90 0.31 38.39 94.57 23.59 0.38
56.23 87.66 17.57 0.24 41.00 91.52 23.14 0.33
59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
Table 4: Impact of pre-training. To validate the effectiveness of pre-training on Nymeria, we show the P&R model’s performance trained from scratch or pre-trained on HumanML3D.
HD-EPIC MoGaze
Pre-train
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
No pre-train 49.76 83.86 22.18 0.30 37.20 87.50 24.11 0.33
HumanML3D 56.23 87.56 18.96 0.28 38.39 93.45 24.85 0.34
Nymeria (Ours) 59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
5.5 Ablation and Analysis
We ablate the proposed P&R model our two largest datasets: HD-EPIC and MoGaze. These cover both estimated motion (using EgoAllo for HD-EPIC) and MoCap data (in MoGaze), to ensure our ablation covers both cases.
Condition Ablation. As explained in Sec. 4.2, the proposed P&R method uses text, initial state (initial pose and velocity), and target location as conditions to generate motion. We ablate the impact of each of these conditions in Tab. 3. Using the object location condition gives a significant boost in all metrics, with maximum gains of $82.6\%$ for reach success and $50.3\%$ in location error on MoGaze. This showcases the difficulty of the task, and that it is not trivial to synthesise P&R motions without knowledge of the target. Using the initial state of the body as a condition helps to improve the priming ability of the generated motion, leading to a gain of $20.3\%$ in prime success on HD-EPIC. We find that having both initial pose and initial velocity as initial state conditions is important, especially for the prime success, with drops of at least $2.3\%$ when either is removed on HD-EPIC. Finally, the ablations show that having action knowledge via text (i.e., drop or pick) also improves P&R motion generation on most metrics.
Impact of pre-training. We find that the Nymeria pre-trained model gives a better initialisation (Tab. 4). This is probably because Nymeria has fine-grained motion for everyday activities; therefore, it infuses the knowledge that humans naturally prime and reach objects for everyday actions, which serves better for the downstream task.
6 Conclusion and Future Work
Humans naturally spot or prime an object before reaching it. Previous motion synthesis benchmarks or methods have failed to explore the role of priming for object reaching. To that end, we curate Prime and Reach (P&R) sequences from five datasets using gaze information and object locations. We propose a P&R motion diffusion model that generates full-body motion using goal pose or target location as a condition, along with initial state and text conditioning. We show that the P&R model can generate better prime and reach motion than previous baselines.
Limitation The current method, in line with similar methods [diomataris2024wandr], does not model hand pose (only body up to wrist). Generating hand motion is an interesting future direction due to its relevance to grasping objects upon reach.
Acknowledgements: This work uses publicly available datasets and annotations to curate P&R sequences. Research at the University of Bristol is supported by EPSRC UMPIRE (EP/T004991/1). M Hatano is supported by JST BOOST, Japan Grant Number JPMJBS2409, and Amano Institute of Technology. S Sinha and J Chalk are supported by EPSRC DTP studentships.
At Keio University, we used ABCI 3.0 provided by AIST and AIST Solutions.
At the University of Bristol, we acknowledge the use of resources provided by the Isambard-AI National AI Research Resource (AIRR), funded by the UK Government’s Department for Science, Innovation and Technology (DSIT) via UK Research and Innovation; and the Science and Technology Facilities Council [ST/AIRR/I-A-I/1023]. In particular, we acknowledge the usage of GPU Node hours granted as part of the Sovereign AI Unit call project “Gen Model in Ego‑sensed World” (Aug-Nov 2025) as well as the usage of GPU Node hours granted by AIRR Early Access Project ANON-BYYG-VXU6-M (March-May 2025).
[…]
Supplementary Material
Appendix A Qualitative Videos
We include the qualitative video on our website https://masashi-hatano.github.io/prime-and-reach/ showcasing predicted motion sequences from our P&R model over different datasets. For each of the sequences, we provide the goal location in green sphere, our goal-pose conditioned prediction in yellow, and the goal-location conditioned synthesis in brown.
Appendix B Further Details on P&R sequence curation
B.1 Slab Test Method for Priming
The Slab Test Method expects the knowledge of the target location $o_{\text{3D}}$, which is an axis-aligned 3D bounding box, defined by its minimum ($\mathbf{b}_{\text{min}}$) and maximum ($\mathbf{b}_{\text{max}}$) corners, or as 3D coordinates of the object center.
The Slab Test Method treats the box as the overlapping volume of three infinite slabs (one for each axis), each bounded by a pair of parallel planes. A visualisation of the intersection checks is shown in Figure S1. The algorithm calculates two key parametric distances along the gaze ray. The first, $t_{\text{near}}$, represents the distance to the last slab plane that the ray enters. It is the furthest entry point, marking the moment the ray is inside all three slabs and thus inside the box. The second, $t_{\text{far}}$, is the distance to the first slab plane that the ray exits. It is the nearest exit point, marking the moment the ray leaves the box volume. A valid intersection occurs if the ray enters the box before it exits, as defined by the condition in Equation 5,
$t_{\text{near}}=\max_{i\in{x,y,z}}\min\left(\frac{\mathbf{b}^{(i)}_{\text{min}}-\mathbf{o}^{(i)}_{\text{cam}}}{\hat{\mathbf{d}}^{(i)}_{\text{gaze}}},\frac{\mathbf{b}^{(i)}_{\text{max}}-\mathbf{o}^{(i)}_{\text{cam}}}{\hat{\mathbf{d}}^{(i)}_{\text{gaze}}}\right)$
$t_{\text{far}}=\min_{i\in{x,y,z}}\max\left(\frac{\mathbf{b}^{(i)}_{\text{min}}-\mathbf{o}^{(i)}_{\text{cam}}}{\hat{\mathbf{d}}^{(i)}_{\text{gaze}}},\frac{\mathbf{b}^{(i)}_{\text{max}}-\mathbf{o}^{(i)}_{\text{cam}}}{\hat{\mathbf{d}}^{(i)}_{\text{gaze}}}\right)$
$\text{Intersection if }t_{\text{near}}<t_{\text{far}}\text{ and }t_{\text{far}}\geq 0,$ (5)
where $\mathbf{o}_{\text{cam}}$ and $\hat{\mathbf{d}}_{\text{gaze}}$ denote the location of the camera and direction of gaze originating from the camera, respectively.
Figure S1: Visualisation of the Slab Test Method [Majercik2018Voxel] for registering primed object interactions. (Left) Primed by Valid Intersection. (Middle) Primed by Near Miss. (Right) No Priming.
To account for near misses where gaze is directed towards an object but does not intersect its bounding box, we employ a proximity check. First, for a given gaze ray originating at $\mathbf{o}_{\text{cam}}$ with direction $\hat{\mathbf{d}}_{\text{gaze}}$, we find the point on the ray, $\mathbf{p}_{\text{closest}}$, that has the minimum distance to the centre of the object’s 3D bounding box, $\mathbf{b}_{\text{centre}}$. This point is found by projecting the vector from the camera to the box centre onto the gaze ray, as shown in Equation 6.
$t_{\text{closest}}=(\mathbf{b}_{\text{centre}}-\mathbf{o}_{\text{cam}})\cdot\hat{\mathbf{d}}_{\text{gaze}}$
$\mathbf{p}_{\text{closest}}=\mathbf{o}_{\text{cam}}+t_{\text{closest}}\cdot\hat{\mathbf{d}}_{\text{gaze}}$ (6)
From this closest point, we cast a new ray directly towards the bounding box centre, $\hat{\mathbf{d}}_{\text{centre}}$, and use the slab test method to identify where this new ray intersects the box. Specifically, we swap $\mathbf{o}_{\text{cam}}$ for $\mathbf{b}_{\text{centre}}$ and $\hat{\mathbf{d}}_{\text{gaze}}$ for $\hat{\mathbf{d}}_{\text{centre}}$ in Equation 5, yielding a point:
$\mathbf{p}_{\text{intersect}}=\mathbf{b}_{\text{centre}}+t_{\text{near}}\cdot\hat{\mathbf{d}}_{\text{centre}}$ (7)
A location is considered primed by a near miss if the Euclidean distance, $\delta$, between $\mathbf{p}_{\text{closest}}$ and $\mathbf{p}_{\text{intersect}}$ is below a threshold $\tau$ of 5 cm. This threshold was determined empirically: we found that smaller values risked undercounting valid gaze interactions due to minor inaccuracies in gaze or object bounding boxes, while larger values began to accept ambiguous cases. Formally, priming by near miss occurs when:
[…]
$\text{Near miss if }\delta\leq\tau\text{ and }t_{\text{closest}}\geq 0$ (8)
The second condition in Equation 8 ensures the closest point lies in front of the camera, confirming the user is looking towards the object.
We exclude interactions involving only minimal movement (<< 20 cm) between the initial pose and goal, as they do not represent meaningful interactions. This filtering process refines the dataset and ensures the quality of P&R sequences so that primed object interactions are not trivial.
B.2 Estimating Full Body Pose for P&R Sequences
Building upon the priming data collected previously, we require full-body pose sequences of primed object interactions. Our generation pipeline uses EgoAllo [yi2025egoallo], a method that estimates expressive, full-body human motion from egocentric video and SLAM-based camera poses. The model first converts head pose trajectories into a spatially and temporally invariant representation that encodes relative motion with respect to the ground plane. This representation is used to condition a diffusion-based prior that samples local SMPL-H [MANO:SIGGRAPHASIA:2017] parameters: pose, representing per-joint rotations over time for the full body including hands; shape, encoding time-invariant body proportions such as height and limb length; and contact predictions, indicating per-joint contact with the environment to improve realism. The model is trained on human motion sequences from AMASS [mahmood2019amass], augmented with synthetic egocentric head pose trajectories.
For each interaction, we provide the model with a sequence of video frames and their corresponding camera poses to generate an initial sequence of full-body motions. To enhance the fidelity of hand-object interactions, we calculate the 3D wrist and palm poses from Aria MPS models and provide these to the EgoAllo model to align the generated hands with the wrist and hand locations. This alignment step is crucial; we found that without it, the hands in the generated sequence often remain static and unrealistic. Incorporating these poses yields a more accurate representation of hand orientation in our final motion sequences.
A key design choice in our generation process is the temporal window of the sequences. Specifically, we initiate the generation 2 seconds prior to the moment the object is primed and conclude following the interaction. This decision was made to ensure that our sequences capture any sufficient head motions or other preparatory body movements that precede the explicit eye-gaze priming. By including this anticipatory phase, the resulting sequences provide a more complete and naturalistic depiction of a primed interaction.
B.3 More Statistics of P&R sequences
We show more detailed statistics on each of our curated datasets in Fig. S2. Concretely, the histograms of body movement, hand movement, and prime gap are shown. Body and hand movement measure the maximum displacement of the body or hands within a P&R motion sequence. Prime gap is the duration between the prime time $t_{p}$ and the pick/put event time $t_{e}$.
Figure S2: Histograms of body movement, hand movement, and prime gap in curated P&R sequences.
B.4 Train-Test Splits
For each dataset, we split the source videos into 70% train -30 % test sets. The P&R sequences curated from these videos were automatically distributed to the corresponding subset. HD-EPIC [perrett2025HD-EPIC] has 156 long videos. We selected 70% (109 videos) for training and the remaining for testing. The curated sequences from the 109 videos were used as train P&R sequences. We perform a similar procedure for MoGaze [kratzer2020mogaze], HOT3D [banerjee2025hot3d], and ADT [pan2023aria]. Zheng et al. [zheng2022gimo] proposed a train-test split for GIMO sequences. We use the same split for our curated P&R sequences. Exact train/test split sizes are given in Tab. S1.
Table S1: Train/Test splits. We provide the train-test splits for our curated P&R sequences.
HD-EPIC MoGaze HOT3D ADT GIMO
Train 12642 1947 1672 326 108
Test 5492 690 744 85 22
Total 18134 2637 2416 411 130
Appendix C Ablation of Architecture & Loss
As in the main paper, our ablations are evaluated on HD-EPIC and MoGaze.
C.1 Transformer Encoder v/s Decoder
We compare performance of transformer encoder v/s decoder based diffusion model for the task of P&R motion generation in Tab. S2. While the decoder architecture injects condition $\mathbf{\hat{z}_{t}}$ by cross-attention with each decoder layer, the encoder provides the condition as an additional token at the input of the first encoder layer. The decoder architecture performs significantly better than the encoder architecture, making it a superior choice for the task.
Table S2: Encoder v/s Decoder. We compare encoder and decoder architecture for P&R motion generation.
HD-EPIC MoGaze
Architecture
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Encoder 51.59 82.91 22.82 0.29 37.95 56.62 27.90 0.34
Decoder 59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
C.2 Training Loss
We ablate the impact of $\mathcal{L}_{joint}$ in Tab. S3. We find adding the $\mathcal{L}_{joint}$ helps improve P&R generation for both HD-EPIC and MoGaze.
Table S3: Loss Ablation. The improvement by adding the joint loss is evident across both datasets and 7 out of 8 metrics.
HD-EPIC MoGaze
Loss
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
$\mathcal{L}$ 57.70 86.70 16.67 0.31 40.10 86.38 21.21 0.34
$\mathcal{L}+\mathcal{L}_{joint}$ 59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
C.3 Incorporating Goal Condition
We pretrain our P&R diffusion model for motion generation only conditioned on text using the Nymeria dataset. For fine-tuning on P&R sequences, we add our initial state and goal pose/target location condition $\mathbf{p}$ to the text condition $\mathbf{z_{t}}$ to get $\mathbf{\hat{z}}_{t}$. We ablate another alternative of incorporating $\mathbf{p}$ to $\mathbf{z_{t}}$ using cross-attention as shown in
$\mathbf{\delta_{t}}=CA(\mathbf{z_{t}},\mathbf{p})=\text{Softmax}\left(\frac{(\mathbf{z_{t}}\mathbf{W}_{Q})(\mathbf{p}\mathbf{W}_{K})^{T})}{\sqrt{d_{k}}}\right)(\mathbf{p}\mathbf{W}_{V})$
$\mathbf{\hat{z}_{t}}=\mathbf{z_{t}}+\mathbf{\delta_{t}}$ (9)
where $CA$ is a 1-layer cross-attention. $\mathbf{z_{t}}$ is linearly projected to get the query and $\mathbf{p}$ is projected to key and value. We use a residual network to make the most of our pretraining. We show the results in Tab. S4. We find that incorporating the condition through addition performs better, across all metrics.
Table S4: Condition Injection. We verify different methods for injecting our initial state and goal conditions.
HD-EPIC MoGaze
Method
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
CACA 56.41 85.67 17.33 0.27 39.43 72.62 26.04 0.35
Addition 59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
C.4 Ablation of Total Diffusion Steps
For a single motion generation, the diffusion model starts from noise at $t=T$ and iteratively denoises it over diffusion steps $t=\{T,T-1,\cdots,0\}$, finally producing the clean motion at $t=0$. We ablate the choice of $T$ in Tab. S5, which controls the total number of steps needed to generate a sequence of motion. We find that $T=50$ gives a consistently good performance across all metrics with an improvement of $1.64\%$ in prime success. Importantly the method is generally robust to the number of steps.
Table S5: Impact of diffusion steps $T$. We compare the performance of P&R motion generation for multiple diffusion steps.
HD-EPIC MoGaze
$T$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
Prime
Success
$\uparrow$
Reach
Success
$\uparrow$
Loc Err $\downarrow$ MPJPE $\downarrow$
10 57.42 90.26 16.00 0.26 41.22 89.88 26.64 0.34
50 59.06 89.48 16.28 0.24 41.44 93.53 22.92 0.33
100 56.18 85.07 17.38 0.27 38.02 91.29 22.25 0.33
500 55.54 85.48 18.13 0.29 36.76 91.96 23.88 0.34
1000 54.68 85.22 17.67 0.29 37.65 90.25 26.04 0.34
Figure S3: Varying time window $\sigma$ and angular error threshold $\theta$ for Prime Success calculation on HD-EPIC and MoGaze.
Appendix D Analysing hyper-parameters of Prime Success metric
We conduct an in-depth analysis to better understand the impact of hyperparameters (the time window $\sigma$ and threshold for angular error $\theta$) used in calculating the newly introduced Prime Success metric. We compare our P&R predictions with the actual gaze while varying the hyperparameters of the metric. Note that as the thresholds are changed, the motion can be considered a success or a failure. Recall that our results are reported for $\theta=16\deg$ and $\sigma=0.2$ sec.
To evaluate the impact of these hyperparameters, we vary $\theta$ on the x-axis (between 0 and 90 $\deg$), then plot distinct curves for discrete time windows: 0, 0.2, 0.4, 0.8, and 1.0 seconds. Figure S3 shows that a very tight time window is too restrictive for MoGaze. As expected, a high $\theta$ threshold is too permissive and cannot be used to compare different methods.
Appendix E Pretraining Results
Here we provide the results of our text-conditioned motion generation pre-training on Nymeria.
As in previous works [tevet2023mdm], we train motion-text embedding models [Guo2022t2m], with two encoders: one for motion and one for text, using a contrastive loss. We use paired text-motion sequences from the Nymeria train set. We follow the architecture for our encoders from  [Guo2022t2m].
Following [guo2020action2motion, tevet2023mdm], we use the following metrics for evaluation -
  • R Precision (Top-3): Given batches of motion and corresponding text, the most similar texts to each motion are ranked based on the Euclidean distances. This calculates the percentage of motion sequences for which the correct text is retrieved in the top 3 matches.
  • FID: This compares the encoded feature distribution of the generated motion to that of real motion
  • Multimodal Distance: This calculates the average Euclidean distance in the embedding space between paired motion and text.
  • Diversity: This measures the variance in the generated motion over all text prompts in the test set.
E.1 Evaluating Embedding and Encoders
We first evaluate the quality of our trained encoders. We use the embedding from [tevet2023mdm] as our baseline, which is trained on HumaML3D. We compare the encoders on the matched text and real motion pairs of the Nymeria test set in Tab. S6.
Table S6: Evaluating encoders.
Feature
Extractor
R Precision
(Top- 3)
$\uparrow$
FID $\downarrow$
Multi-modal
Distance
$\downarrow$
Diversity $\uparrow$
[…]
Different from HumanML3D, Nymeria text and motion relate to more fine-grained everyday activities e.g., ‘In the hallway, C slightly turns her body to the left as she looks at her peer while holding the right side door of the laundry closet using her right hand. Then C slightly turns her body to the right as she looks at the laundry closet and takes her right hand off from the right side door of the laundry closet’. As a result, HumanML3D-trained text embeddings do not work well on the Nymeria test set. The high R-Precision and low multi-modal distance show that our trained embedding on Nymeria better matches text-motion pairs while maintaining high diversity. This shows that our trained embedding is substantially better and we thus use it for evaluating the impact of our pretraining.
E.2 Pretraining Results
We provide the results of our pretrained model in Tab. S7. We find that the motion generated by our Nymeria-pretrained model aligns better with the fine-grained texts of Nymeria. This is verified by the $34.0\%$ and 2.68-2.68 improvements in R-Precision and Multi-modal distance respectively. The diversity of our generated motions is +2.49+2.49 higher than that of motions generated by the HumanML3D pre-trained model. We showcase some of the qualitative results of the pre-trained model in Fig. S4.
Table S7: Pretraining results.
Pretraining
Dataset
Motion
R Precision
(Top- 3)
$\uparrow$
FID $\downarrow$
Multi-modal
Distance
$\downarrow$
Diversity $\uparrow$
[…]
Figure S4: Qualitatives of our pre-trained model. We showcase both ground truth (GT) and predicted motion for given text prompts. Darker poses represent later times.